# Multi-aspect Knowledge Distillation with Large Language Model

### v3

@@ 4.3 Ablation studies
Effect of the loss function. In Table 3 (a), we investigate the effect of the loss function by applying KL-divergence loss to the multi-aspect logit. The result shows that using binary-cross entropy loss achieves better performance. We assume that because the multi-aspect logits represent the probability of the yes token extracted from the MLLM, using binary-cross entropy loss would bring more improvement to the classification model.
Effect of the multi-aspect logits. In Table 3 (b), we validate the contribution of the multi-aspect logits to image classification by comparing our method to the one that replaces the logits with a random logit following a Gaussian distribution. As shown in Table 3 (b), our method with multi-aspect logits outperforms the method with random logits. These results demonstrate that the multi-aspect logits can enhance image classification performance by representing knowledge from various aspects for each class in the dataset.
Weight to the multi-aspect knowledge distillation loss. Table 3 (c) presents the performance of our method with different weights to the multi-aspect logit loss on StanfordCars and Caltech101. The x-axis represents the weights $\alpha$ (0 means the baselines), while the y-axis indicates the accuracy. Our method, based on $\alpha$, demonstrates improvements in the performances of all baseline models. Additionally, we empirically find that the performance decreases when $\alpha$ value reaches 50.
Effect of LLM on multi-aspect question generation. To assess the impact of different LLMs on multi-aspect question generation, we compare a model that generates multi-aspect questions using GPT-3.5 with our model that generates multi-aspect questions using GPT-4o. Both models utilize InternVL2-8B as the MLLM for logit extraction, with only the LLM for multi-aspect question generation being different. In Table 3 (d), Ours(L:GPT-3.5) using GPT-3.5 for generating multi-aspect questions outperforms the baselines and shows competitive results when compared to ours(which uses GPT-4o). These results demonstrate the robustness of our method to the performance of LLMs.
Effect of MLLM on multi-aspect logit extraction. We further investigate the impact of using different MLLMs on our method by using LLaVA-NeXT-34B Liu et al. (2024a), which has more parameters compared to InternVL2-8B Chen et al. (2024). As shown in Table 3 (d) with Ours (M: LLaVA), our method with LLaVA-NeXT-34B outperforms the baselines and shows competitive results when compared to InternVL2-8B. However, InternVL2-8B is more parameter efficient.
Figure 2: Ablation study on the number of multi-aspect questions. The x-axis represents the number of aspects (0 means the baselines), while the y-axis indicates the accuracy. We run each experiment three times and report the average results.
Table 3: Ablation study on each component. Table (a), (b) and (d) report the accuracy (%) on StanfordCars Krause et al. (2013). Figure (c) shows different weights to the multi-aspect loss on StanfordCars and Caltech101. Res18 for ResNet18, Res34 for ResNet34, Mb-N1 for MobileNetV1 and EffiNet for EfficientNet. Rand for our method with random logits instead of multi-aspect logits. KL for our method with KL-Divergence loss on multi-aspect logit. $\alpha$ for the weighting factor of multi-aspect logit loss. We run each experiment three times and report the average results. We provide additional experimental results in the supplementary material.
(a) Effect of the loss function
@@
Ours 83.38 84.33 85.43 88.07
(b) Effect of the multi-aspect logit
Res18 Res34 Mb-N1 EffiNet
@@
(c) Weights to the multi-aspect loss
(d) Effect of LLM and MLLM
@@
Teacher MLLM (85.52)
Student Res18 Res34
Base 73.35 75.36
KD 73.86 75.86
Ours 75.76 77.56
Table 5: Extension to less training data. Data represents the percentage of training data used, while the Gap indicates the gap in accuracy between the baseline and our method with ResNet18. Base is the baseline using cross-entropy loss with class labels.
@@
Table 6: Extension to traditional knowledge distillation on StanfordCars and Caltech101. We run each experiment three times and report the average results.
Teacher Res34(80.93) EffiNet(86.41)
Dataset Student Res18(77.53) Mb-N1(82.84)
Stanford Cars KD 79.62 85.11
DKD 82.55 85.93
Ours + KD 83.44 86.34
Ours + DKD 83.23 86.43
Teacher Res34(75.36) EffiNet(80.05)
Dataset Student Res18(73.35) Mb-N1(76.64)
Caltech 101 KD 74.53 78.71
DKD 76.37 79.95
Ours + KD 76.70 79.70
Ours + DKD 77.41 80.95
Table 7: Extension to object detection on MS-COCO based on Faster-RCNN Ren et al. (2016)-FPN Lin et al. (2017). AP evaluated on val2017. We run each experiment three times and report the average results.
@@
4.4 Extension of our model
To show the scalability of our approach, we apply our method to three tasks. First, we extend our model using traditional logit distillation. Second, we evaluate our model’s performance when the dataset size is decreased. Finally, we extend our model to the object detection task.
Extension to traditional knowledge distillation. Since our model does not have the teacher classification model and the teacher model’s class logits, it is different from traditional knowledge distillation (KD). However, since we distill the multi-aspect knowledge to be learned into logits, it simply can be integrated with existing logit distillation methods. We compare our method with KD and DKD Zhao et al. (2022) on the StanfordCars Krause et al. (2013) and Caltech101 Fei-Fei et al. (2004). According to Table 4.3, extending our approach to traditional KD approaches outperforms traditional KD approaches. These results demonstrate that our approach can be effectively extended to both traditional logit distillation and DKD method.
Extension to less training data. We evaluate the performance of our model when trained with a reduced amount of training data. As shown in Table 5, our multi-aspect approach leads to greater performance improvement as the dataset size decreases. For example, on the StanfordCars dataset, ResNet18 shows a 24.01% performance improvement over the baseline when only 40% of the entire training dataset was used. It demonstrates the potential for broader applicability in fine-grained tasks and real-world applications with limited training datasets.
Extension to object detection. To evaluate the scalability of our method, we evaluate the performance on object detection tasks with MS-COCO datasets. Following Zhao et al. (2022), we add features to the backbone network of Faster R-CNN Ren et al. (2016)-FPN Lin et al. (2017) and apply a multi-aspect logit loss with the number of multi-aspect questions set to 50. As shown in Table 7, our method further improves the performances of the baselines. These results show that we can effectively identifying objects in the image by learning deep visual feature from multi-aspect knowledge and may have a potential to contribute to various visual understanding tasks.
Extension to large parameter models. While we focus on challenging yet practical fine-grained datasets and small models, we also evaluate scalability to larger models through experiments on large-parameter models such as ResNet101 and ViT models. According to Table 5.2, our model outperforms the baselines with larger parameter models such as ResNet101 and ViT. These results demonstrate that our model can be effectively scaled to pre-trained and large parameter models.
@@ 5.1 Distillation with MLLM zero-shot classification logits
According to Table 1, the MLLM shows poor zero-shot image classification performance on fine-grained datasets. These results show that they may struggle with classifying highly specific information, such as distinguishing between Yellow headed Blackbird and Eastern Towhee in the CUB200 Wah et al. (2011) dataset. Therefore, we cannot directly distill the class logits from MLLM. To leverage the features of MLLM that can understand and infer abstract and complex information, we distill knowledge through multi-aspect questions based on diverse insights and understanding beyond class labels. This shows the potential of our approach to be applied to other tasks, regardless of the performance of MLLM in specific domains.
In coarse-grained image datasets, we find that MLLM performs better than on fine-grained datasets. We assume that this is because MLLM was trained on a very large dataset, enabling it to perform general classification tasks. Since the zero-shot classification performance of MLLM on Caltech101 is better than the baseline, we may apply traditional knowledge distillation (KD) using MLLM’s class logits as the teacher logits on Caltech101. According to Table 4.3, using MLLM’s logits as a teacher result in a slight performance improvement over the baseline, but it underperforms compared to our method. Additionally, when applying our approach to coarse-grained image dataset, it improve the performance of all models over the baselines, as shown in Table 2. This shows that not only for fine-grained but also for coarse-grained tasks, it is important to consider multi-aspects rather than directly distilling the logits of MLLM, demonstrating that our approach is more effective.
Figure 3: Visualization of the average logit distribution for classes related to aspects. The x-axis represents the classes, and the y-axis represents the mean of the aspect probability distribution from MLLMs in the dataset. The class names corresponding to the indices in x-axis are provided in the supplementary material due to space.
Figure 4: Visualization of t-SNE embeddings for the datasets by aspects. Ours is t-SNE visualizations of the aspect logits from our model (ResNet18), while MLLM is t-SNE visualizations of the aspect logits from the MLLM (InternVL2-8B). The yellow points indicate that the probability of “yes” is close to 1, and the purple points indicate that the probability of “yes” is close to 0.
5.2 Analysis of multi-aspect questions generated by the LLM
To analyze the effectiveness of the multi-aspect questions generated by the LLM in image classification, we present a histogram of the average MLLM probability values of aspects for each class, as shown in Figure 3. For example, as shown in Figure 3 (a)-1, the class “BMW M6 Convertible 2010” on StanfordCars Krause et al. (2013) has a high probability value for the aspect “Does the car have a convertible roof?”. We observe that classes possessing the features of the aspect exhibit high probabilities, while those lacking the features show low probabilities.
@@ 5.2 Analysis of multi-aspect questions generated by the LLM
Table 8: Extension to large parameter models on StanfordCars and Mini-ImageNet. * indicates that the model was trained using a pretrained model from ImageNet-1K. We run each experiment three times and report the average results.
StanfordCars Mini-ImageNet
Model Base Ours Gap Base Ours Gap
ResNet101 81.36 83.97 +2.62 77.00 78.11 +1.11
ViT-B/16 10.88 11.62 +0.74 43.41 44.18 +0.77
ViT-B/16* 85.72 86.13 +0.41 96.64 96.87 +0.23
ViT-B/32 9.32 9.73 +0.41 41.42 42.73 +1.31
ViT-B/32* 79.00 79.98 +0.98 93.41 93.50 +0.09
Figure 5: Comparison of probability values for multi-aspect questions. We compare the probability values of our model with those of the MLLM for multi-aspect questions. Our model shows similar probability values to MLLM across various multi-aspect questions.
5.3 Analysis of the discriminability using the aspect logits
@@
5.4 Analysis of multi-aspect classification of our model
To analyze the classification performance of our model for multi-aspect questions, we compare the probability values of our model with those of the MLLM for multi-aspect questions. As shown in Figure 5 (c), when an image of a Birman is given as input, our model outputs a probability value of 86.97 for the visual aspect “Does the animal have striking blue eyes?” and a value of 11.74 for the aspect “Does the animal have floppy ears?”, similar to the MLLM. These results indicate that our model effectively distill visual aspects and understands visual aspects.
Furthermore, as shown in Figure 5 (d), when an image of a Leopards is given as input, our model outputs a probability value of 96.23 for the aspect “Is the object known for its speed or ability to move quickly?” and a value of 98.46 for the aspect “Is the object typically found outdoors in a natural environment?” which are not visual aspect but abstract or require a deeper understanding of the image, similar to the MLLM. These results suggest that the model can distill not only visual knowledge but also abstract and complex knowledge about multi-aspect knowledge.
As shown in Figure 6, we provide two types of training loss curves with the number of iterations: (a) cross-entropy loss and (b) our proposed MaKD loss. When $\alpha=0$, it represents without MaKD loss, which means the baseline model. (a) shows the cross-entropy loss curve according to the $\alpha$ values of MaKD, while (b) presents the normalized MaKD loss curve according to $\alpha$ values without cross-entropy loss. According to (a), applying our method results in faster convergence in training and the model learns multi-aspect effectively with decreasing MaKD loss. These results demonstrate that our method effectively learns multi-aspect knowledge.
Figure 6: The training curve graph of loss with the number of iterations on StanfordCars. We provide two types of training curve losses. (a) cross-entropy loss and (b) our proposed MaKD loss. Applying our method results in faster convergence in training and the model learns multi-aspect effectively with decreasing MaKD loss.
@@ 5.5 Training time and computational cost
As we extract logits from MLLMs, this can require more computational resources compared to training only image classification models. However, since we query the MLLM about aspects in a zero-shot manner, there is no need to train the MLLM. Moreover, we utilize InternVL2-8B Chen et al. (2024) for logit extraction, which allows aspect extraction using a single NVIDIA RTX 3090. The number of parameters in our model is approximately 11.25M when using ResNet18 with 50 aspects, with the baseline also having 11.23M parameters. For StanfordCars, the training time for the baseline model is 25.42 seconds per epoch, while our model takes 27.90 seconds per epoch. In terms of inference time, our model takes 22.80 seconds, compared to the baseline’s 20.59 seconds, showing slight increase. Also we provide the time required for MLLMs to extract aspect responses from the training dataset. For the StanfordCars dataset, it takes approximately 0.83 seconds ± 174 ms per image for a single aspect, while for Mini-ImageNet, it takes about 0.942 seconds ± 169 ms. The annotation time varies depending on the number of aspects and the size of the dataset.
@@ 6 Conclusion and Limitation
In this paper, we propose a novel multi-aspect knowledge distillation method leveraging MLLM along with analyses. Unlike previous image classification methods, our method leverages MLLM to distill multi-aspect knowledge that require complex and deeper understanding beyond the class labels. Our experimental results demonstrate that the proposed method outperforms baseline models in both fine-grained and course-grained image classification tasks. Additionally, we extend our method to other tasks such as object detection, and it outperforms the baselines. Our findings provide a novel view by simply distilling multi-aspect knowledge and demonstrate the potential of our method to be applied to a variety of tasks. However, as a limitation, our approach is constrained by the necessity of pre-trained LLMs and MLLMs to generate aspects and logits used for model training. In future work, we will explore applying our method to other domains, such as image generation and image captioning.

### v4

@@ 4.3 Ablation Studies
Effect of the loss function. In Table 2, we investigate the effect of the loss function by applying KL-divergence loss to the multi-aspect logit. The result shows that using binary-cross entropy loss achieves better performance. We assume that because the multi-aspect logits represent the probability of the yes token extracted from the MLLM, using binary-cross entropy loss would bring more improvement to the classification model.
@@
Table 2: Ablation study on multi-aspect logits and loss function. We report the accuracy (%) on StanfordCars  [10] dataset. KL for our method with KL-Divergence loss on multi-aspect logit. Rand for our method with random logits instead of multi-aspect logits.
@@
Table 3: Ablation study on multi-aspect question generation from LLM and logit extraction from MLLM. We report the accuracy (%) on StanfordCars  [10] dataset. L means that we used GPT-3.5 as the LLM for generating multi-aspect questions, while M indicates that we used LLaVA as the MLLM for extracting multi-aspect logits.
Figure 2: Ablation study on the number of multi-aspect questions. The x-axis represents the number of aspects (0 represents the baseline model), while the y-axis indicates the accuracy. We run each experiment three times and report the average results.
Effect of the multi-aspect logits. In Table 2, we validate the contribution of the multi-aspect logits to image classification by comparing our method to the one that replaces the logits with a random logit following a Gaussian distribution. As shown in Table 2, our method with multi-aspect logits outperforms the method with random logits. These results demonstrate that the multi-aspect logits can enhance image classification performance by representing knowledge from various aspects for each class in the dataset.
Weight to the multi-aspect knowledge distillation loss. Figure 3 presents the performance of our method with different weights to the multi-aspect logit loss on StanfordCars and Caltech101. The x-axis represents the weights $\alpha$ (0 means the baselines), while the y-axis indicates the accuracy. Our method, based on $\alpha$, demonstrates improvements in the performances of all baseline models. Additionally, we empirically find that the performance decreases when $\alpha$ value reaches 50.
Figure 3: Ablation study on weight to the multi-aspect knowledge distillation loss. $\alpha$ value is the weighting factor of multi-aspect logit loss.
Effect of LLM on multi-aspect question generation. To assess the impact of different LLMs on multi-aspect question generation, we compare a model that generates multi-aspect questions using GPT-3.5 with our model that generates multi-aspect questions using GPT-4o. Both models utilize InternVL2-8B as the MLLM for logit extraction, with only the LLM for multi-aspect question generation being different. In Table 3, Ours (L: GPT-3.5) using GPT-3.5 for generating multi-aspect questions outperforms the baselines and shows competitive results when compared to ours (which uses GPT-4o). These results demonstrate the robustness of our method to the performance of LLMs.
Effect of MLLM on multi-aspect logit extraction. We further investigate the impact of using different MLLMs on our method by using LLaVA-NeXT-34B [13], which has more parameters compared to InternVL2-8B [4]. As shown in Table 3 with Ours (M: LLaVA), our method with LLaVA-NeXT-34B outperforms the baselines and shows competitive results when compared to InternVL2-8B. However, InternVL2-8B is more parameter efficient.
4.4 Extension of Our Model
To show the scalability of our approach, we first extend our model using traditional logit distillation. Second, we evaluate our model’s performance when the dataset size is decreased. Finally, we apply our model to object detection.
Extension to traditional knowledge distillation. Since our model does not have the teacher classification model and the teacher model’s class logits, it is different from traditional KD. However, since we distill the multi-aspect knowledge to be learned into logits, it simply can be integrated with existing logit distillation methods. We compare our method with KD and Decoupled Knowledge Distillation (DKD) [43] on the StanfordCars [10] and Caltech101 [6]. According to Table 4, extending our approach to traditional KD approaches outperforms traditional KD approaches. These results demonstrate that our approach can be effectively extended to both traditional KD and DKD.
Res34 (80.93) EffiNet (86.41)
Dataset / Student Res18 (77.53) Mb-N1 (82.84)
StanfordCars KD 79.62 85.11
DKD 82.55 85.93
Ours + KD 83.44 86.34
Ours + DKD 83.23 86.43
Res34 (75.36) EffiNet (80.05)
Dataset / Student Res18 (73.35) Mb-N1 (76.64)
Caltech101 KD 74.53 78.71
DKD 76.37 79.95
Ours + KD 76.70 79.70
Ours + DKD 77.41 80.95
Table 4: Extension to traditional knowledge distillation on StanfordCars and Caltech101. KD is traditional knowledge distillation and DKD is decoupled knowledge distillation [43]. We run each experiment three times and report the average results.
@@ 4.4 Extension of Our Model
Table 5: Extension to class logit distillation with MLLM on Caltech101. Base is the baseline using cross-entropy loss with class labels. We run each experiment three times and report the average results.
@@
Table 6: Extension to object detection on MS-COCO based on Faster-RCNN [25]-FPN [12]. AP evaluated on val2017. We run each experiment three times and report the average results.
Extension to less training data. We evaluate the performance of our model when trained with a reduced amount of training data. As shown in Table 5, our multi-aspect approach leads to greater performance improvement as the dataset size decreases. For example, on the StanfordCars dataset, ResNet18 shows a 24.01% performance improvement over the baseline when only 40% of the training dataset was used. It demonstrates the potential for broader applicability in fine-grained tasks and real-world applications with limited training datasets.
Application to object detection. We apply our method to object detection with MS-COCO datasets. Following [43], we add features to the backbone network of Faster R-CNN [25]-FPN [12] and apply a multi-aspect logit loss with the number of multi-aspect questions set to 50. As shown in Table 6, our method demonstrates a trend of marginal improvement over the baselines. It shows potential for other visual understanding tasks.
Figure 4: Visualization of the average logit distribution for classes related to aspects. The x-axis represents the classes, and the y-axis represents the mean of the aspect probability distribution from MLLM in the dataset.
Figure 5: Visualization of t-SNE embeddings for the datasets by aspects. Ours is t-SNE visualizations of the aspect logits from our model (ResNet18), while MLLM is t-SNE visualizations of the aspect logits from the MLLM (InternVL2-8B). The yellow points indicate that the probability of “yes” is close to 1, and the purple points indicate that the probability of “yes” is close to 0. Our model exhibit a similar trend to the aspect logits of the MLLM in fine-grained datasets  [22, 19].
@@ 5.1 Distillation with MLLM Zero-shot Classification Logits
According to Table 1, the MLLM shows poor zero-shot image classification performance on fine-grained datasets. These results show that they may struggle with classifying highly specific information, such as distinguishing between Yellow-headed Blackbird and Eastern Towhee in the CUB200 [32] dataset. Therefore, we cannot directly distill the class logits from MLLM. To leverage the features of MLLM that can understand and infer abstract and complex information, we distill knowledge through multi-aspect questions based on diverse insights and understanding beyond class labels. This shows the potential of our approach to be applied to other tasks, regardless of the performance of MLLM in specific domains.
In coarse-grained image datasets, we find that MLLM performs better than on fine-grained datasets. We assume that this is because MLLM was trained on a very large dataset, enabling it to perform general classification tasks. Since the zero-shot classification performance of MLLM on Caltech101 is better than the baseline, we may apply traditional KD using MLLM’s class logits as the teacher logits on Caltech101. According to Table 7, using MLLM’s logits as a teacher result in a slight performance improvement over the baseline, but it underperforms compared to our method. Additionally, when applying our approach to coarse-grained image datasets, it improves the performance of all models over the baselines, as shown in Table 1. This shows that not only for fine-grained but also for coarse-grained tasks, it is important to consider multi-aspects rather than directly distilling the logits of MLLM, demonstrating that our approach is more effective.
Figure 6: Comparison of probability values for multi-aspect questions. We compare the probability values of our model with those of the MLLM for multi-aspect questions. Our model shows similar probability values to MLLM across various multi-aspect questions on test datasets.
Teacher MLLM (85.52)
Student ResNet18 ResNet34
Base 73.35 75.36
KD 73.86 75.86
Ours 75.76 77.56
5.2 Multi-aspect Questions Generated by LLM
To analyze the effectiveness of the multi-aspect questions generated by the LLM in image classification, we present a histogram of the average MLLM probability values of aspects for each class, as shown in Figure 4. For example, as shown in Figure 4 (a)-1, the class “BMW M6 Convertible 2010” on StanfordCars [10] has a high probability value for the aspect “Does the car have a convertible roof?”. We observe that classes possessing the features of the aspect exhibit high probabilities, while those lacking the features show low probabilities.
@@ 5.2 Multi-aspect Questions Generated by LLM
5.3 Discriminability using Multi-aspect Logits
@@
5.4 Multi-aspect Probability Values of Our Model
To analyze the classification performance of our model for multi-aspect questions, we compare the probability values of our model with those of the MLLM for multi-aspect questions. As shown in Figure 6 (c), when an image of a “Birman” is given as input, our model outputs a probability value of 86.97 for the visual aspect “Does the animal have striking blue eyes?” and a value of 11.74 for the aspect “Does the animal have floppy ears?”, similar to the MLLM. These results indicate that our model effectively distill visual aspects and understands visual aspects.
Furthermore, as shown in Figure 6 (d), when an image of a “Leopards” is given as input, our model outputs a probability value of 96.23 for the aspect “Is the object known for its speed or ability to move quickly?” and a value of 98.46 for the aspect “Is the object typically found outdoors in a natural environment?” which are not visual aspects but abstract or require a deeper understanding of the image, similar to the MLLM. These results suggest that the model can distill not only visual knowledge but also abstract and complex knowledge about multi-aspect knowledge.
@@ 5.5 Training Time and Computational Cost
As we extract logits from MLLMs, this can require more computational resources compared to training only image classification models. However, since we query the MLLM about aspects in a zero-shot manner, there is no need to train the MLLM. Moreover, we utilize InternVL2-8B [4] for logit extraction, which allows aspect extraction using a single NVIDIA RTX 3090. The number of parameters in our model is approximately 11.25M when using ResNet18 with 50 aspects, with the baseline also having 11.23M parameters. For StanfordCars, the training time for the baseline model is 25.42 seconds per epoch, while our model takes 27.90 seconds per epoch. Also, we provide the time required for MLLMs to extract aspect responses from the training dataset. For the StanfordCars dataset, it takes approximately 0.83 seconds per image for a single aspect, while for Mini-ImageNet, it takes about 0.942 seconds. The annotation time varies depending on the number of aspects and the size of the dataset.
@@ 6 Conclusion and Limitation
In this paper, we propose a novel multi-aspect knowledge distillation method leveraging MLLM along with analyses. Unlike previous image classification methods, our method leverages MLLM to distill multi-aspect knowledge that requires complex and deeper understanding beyond the class labels. Our experimental results demonstrate that the proposed method outperforms baseline models in both fine-grained and coarse-grained image classification tasks. Our findings offer a novel perspective by simply distilling multi-aspect knowledge and demonstrate the potential of our method to be applied to a variety of tasks.
Limitation and future work. Our approach is constrained by the necessity of pre-trained LLMs and MLLMs to generate aspects and logits used for model training. Furthermore, the performance may be affected by the multi-aspect questions in MLLMs, making generalization across datasets challenging. In future work, we plan to explore optimization strategies that facilitate better generalization of our method, and apply it to other domains such as image captioning and image generation.